# Differential Privacy as a Perk: Federated Learning over Multiple-Access Fading Channels with a Multi-Antenna Base Station

### v2

@@ IV-B Problem Formulation
$(\text{P0}):$ $\min_{\{\boldsymbol{w}^{(t)}\}_{t=0}^{T-1},\ \{s_{i}^{(t)}\}_{t=0}^{T-1}}\ \sum_{t=0}^{T-1}\left(\Lambda^{(t)}+\|\boldsymbol{w}^{(t)}\|^{2}d\sigma^{2}\right)$ (30)
$\text{ s.t. }\quad c^{2}(s_{i}^{(t)})^{2}\leq dP,\ \forall i\in\mathcal{I}^{(t)},\ \forall t,$
$\epsilon\geq\sqrt{\frac{(2c_{\delta}+8)\log(1/\delta)rc^{2}}{\sigma^{2}}\min\left\{\sum_{t=0}^{T-1}\phi_{t},\Phi\right\}}.$
Given any $\{\boldsymbol{w}^{(t)}\}_{t=0}^{T-1}$, the optimal $s_{i}^{(t)}$ is given by [37, 3]
$s_{i}^{(t)}=\frac{1}{(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}},$ (31)
which yields $\Lambda^{(t)}=0$, and $\phi_{t}=1/\|\boldsymbol{w}^{(t)}\|^{2}$. By substituting $s_{i}^{(t)}$, $i\in\mathcal{I}^{(t)}$, $\forall t$, we recast problem (P0) into the following equivalent problem.
$(\text{P0}^{\prime}):$ $\min_{\{\boldsymbol{w}^{(t)}\}_{t=0}^{T-1}}\ \sum_{t=0}^{T-1}\|\boldsymbol{w}^{(t)}\|^{2}$ (32)
$\text{ s.t. }\quad\left|(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}\right|\geq\frac{c}{dP},\ \forall i\in\mathcal{I}^{(t)},\ \forall t,$
$\epsilon\geq\sqrt{\frac{(2c_{\delta}+8)\log(1/\delta)rc^{2}}{\sigma^{2}}\min\left\{\sum_{t=0}^{T-1}\frac{1}{\|\boldsymbol{w}^{(t)}\|^{2}},\Phi\right\}}.$
For illustration, we solve a simplified version of $(\text{P0}^{\prime})$ assuming an unbounded parameter domain (i.e., $D\to\infty$), yielding $\min\{\sum_{t}1/\|\boldsymbol{w}^{(t)}\|^{2},\Phi\}=\sum_{t}1/\|\boldsymbol{w}^{(t)}\|^{2}$. Note that the general case with finite $D$ is also tractable and discussed in Sec. VI. In this case, $(\text{P0}^{\prime})$ is simplified as:
$(\widetilde{\text{P}}\text{0}):\$$\min_{\{\boldsymbol{w}^{(t)}\}_{t=0}^{T-1}}\quad\sum_{t=0}^{T-1}\|\boldsymbol{w}^{(t)}\|^{2}$ (33a)
$\text{s.t.}\quad|(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}|\geq c\mathord{\left/\right.\kern-1.2pt}\sqrt{dP},\ \forall i\in\mathcal{I}^{(t)},\ \forall t,$ (33b)
$\quad\ \ \ \ \sum_{t=0}^{T-1}\frac{1}{\|\boldsymbol{w}^{(t)}\|^{2}}\leq A,$ (33c)
@@
Theorem IV.1
@@
$(\widetilde{\text{P}}\text{0}^{\prime}):\$$\min_{\{\boldsymbol{w}^{(t)}\}_{t=0}^{T-1}}\quad\sum_{t=0}^{T-1}\|\boldsymbol{w}^{(t)}\|^{2}$ (35a)
$\text{ s.t. }\quad\|\boldsymbol{w}^{(t)}\|\geq\|\boldsymbol{w}_{\text{ZF}}^{(t)}\|>0,\ \forall t,$ (35b)
$\quad\ \ \ \ \sum_{t=0}^{T-1}\frac{1}{\|\boldsymbol{w}^{(t)}\|^{2}}\leq A,$ (35c)
with
$\boldsymbol{w}^{(t)}_{\text{ZF}}=\frac{c}{\sqrt{dP}}\boldsymbol{H}^{(t)}\left((\boldsymbol{H}^{(t)})^{H}\boldsymbol{H}^{(t)}\right)^{-1}\boldsymbol{u},$ (36)
where $\boldsymbol{H}^{(t)}=[\boldsymbol{h}_{1}^{(t)},\cdots,\boldsymbol{h}_{rn}^{(t)}]\in\mathbb{C}^{m\times rn}$, and $\boldsymbol{u}\in\mathbb{C}^{rn}$ denotes the vector whose components all have unit modulus.
Proof: To prove this result, we first establish the optimal solution in the absence of the DP constraint (33c). Without the DP constraint (33c), $(\widetilde{\text{P}}\text{0})$ can be decoupled into $T$ quadratically constrained quadratic programs (QCQPs), each for one $t$ as follows:
$(\text{P1}):\$$\min_{\boldsymbol{w}^{(t)}}\quad\|\boldsymbol{w}^{(t)}\|^{2}$ (37a)
$\text{s.t.}\quad|(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}|\geq c\mathord{\left/\right.\kern-1.2pt}\sqrt{dP},\ \forall i\in\mathcal{I}^{(t)},\ \forall t.$ (37b)
Generally, each sub-problem can be individually addressed by using the semidefinite relaxation (SDR) method [38], which involves relaxing the non-convex rank-$1$ constraint to form a solvable semidefinite program (SDP) [39] and then recovering a rank-$1$ solution. Here, we provide a more direct analytical solution in the following lemma.
Lemma IV.2
Assuming that $m\geq rn$ and i.i.d. channels coefficients, the optimal solution of $(\text{P1})$ is the zero-forcing (ZF) combiner [40]:
$\boldsymbol{w}^{(t)}_{\text{ZF}}=\frac{c}{\sqrt{dP}}\boldsymbol{H}^{(t)}\left((\boldsymbol{H}^{(t)})^{H}\boldsymbol{H}^{(t)}\right)^{-1}\boldsymbol{u},$ (38)
where $\boldsymbol{H}^{(t)}=[\boldsymbol{h}_{1}^{(t)},\cdots,\boldsymbol{h}_{rn}^{(t)}]\in\mathbb{C}^{m\times rn}$, and $\boldsymbol{u}\in\mathbb{C}^{rn}$ denotes the vector whose components all have unit modulus.
The detailed proof of Lemma IV.2 is provided in Appendix F. We now prove that solving $(\widetilde{\text{P}}\text{0})$ and $(\widetilde{\text{P}}\text{0}^{\prime})$ are equivalent.
Denote the optimal solutions of $(\widetilde{\text{P}}\text{0})$ and $(\widetilde{\text{P}}\text{0}^{\prime})$ as $\{\hat{\boldsymbol{w}}^{(t)}\}$ and $\{\widetilde{\boldsymbol{w}}^{(t)}\}$, respectively. First, we note that the solution $\{\hat{\boldsymbol{w}}^{(t)}\}$ also satisfies the constraint of (35b) and (35c) by Lemma IV.2, making it a feasible solution for $(\widetilde{\text{P}}\text{0}^{\prime})$. We then have
@@
In addition, for the solution $\{\frac{\|\widetilde{\boldsymbol{w}}^{(t)}\|}{\|\boldsymbol{w}^{(t)}_{\text{ZF}}\|}\boldsymbol{w}_{\text{ZF}}^{(t)}\}_{t=0}^{T-1}$, we have
$|\left(\frac{\|\widetilde{\boldsymbol{w}}^{(t)}\|}{\|\boldsymbol{w}^{(t)}_{\text{ZF}}\|}\boldsymbol{w}_{\text{ZF}}^{(t)}\right)^{H}\boldsymbol{h}_{i}^{(t)}|$ $\geq|(\boldsymbol{w}_{\text{ZF}}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}|$ (40)
@@
$\sum_{t=0}^{T-1}\frac{1}{\|\frac{\|\widetilde{\boldsymbol{w}}^{(t)}\|}{\|\boldsymbol{w}^{(t)}_{\text{ZF}}\|}\boldsymbol{w}_{\text{ZF}}^{(t)}\|^{2}}$ $\leq\sum_{t=0}^{T-1}\frac{1}{\|\widetilde{\boldsymbol{w}}^{(t)}\|^{2}}$ (41)
$\leq A,$
which means that $\bigl\{\frac{\|\widetilde{\boldsymbol{w}}^{(t)}\|}{\|\boldsymbol{w}^{(t)}_{\text{ZF}}\|}\boldsymbol{w}_{\text{ZF}}^{(t)}\bigr\}_{t=0}^{T-1}$ is a feasible solution for $(\widetilde{\text{P}}\text{0})$. Further, we have
$\sum_{t=0}^{T-1}\|\hat{\boldsymbol{w}}^{(t)}\|^{2}\leq\sum_{t=0}^{T-1}\|\frac{\|\widetilde{\boldsymbol{w}}^{(t)}\|}{\|\boldsymbol{w}^{(t)}_{\text{ZF}}\|}\boldsymbol{w}_{\text{ZF}}^{(t)}\|^{2}\leq\sum_{t=0}^{T-1}\|\widetilde{\boldsymbol{w}}^{(t)}\|^{2}.$ (42)
@@
IV-C Optimal Solution
In this subsection, we proceed to derive the optimal solution of problem $(\widetilde{\text{P}}\text{0}^{\prime})$. Let $\pi_{t}=\|\boldsymbol{w}_{\text{ZF}}^{(t)}\|$ and define $q_{t}=\|\boldsymbol{w}^{(t)}\|$. Then, we can rewrite $(\widetilde{\text{P}}\text{0}^{\prime})$ as
$(\widetilde{\text{P}}\text{0}^{\prime\prime}):\$$\min_{\{q_{t}\}_{t=0}^{T-1}}\quad\sum_{t=1}^{T-1}q_{t}^{2}$ (43a)
$\text{ s.t. }\quad q_{t}\geq\pi_{t}>0,\ \forall t,$ (43b)
$\quad\ \ \ \ \ \ \ \sum_{t=0}^{T-1}\frac{1}{q_{t}^{2}}\leq A,$ (43c)
@@ IV-C Optimal Solution
where $\mu\geq 0$ and $\nu_{t}\geq 0$ are the Lagrangian multipliers associated with constraints (43b) and (43c), respectively. Further, the KKT conditions are given as
@@
where $\mu^{*}$ denotes the optimal value of the dual variable.
@@
Case (i) If $\sum_{t=0}^{T-1}1/\pi_{t}^{2}\leq A$, the solution $\{\pi_{t}\}$ satisfies privacy and is optimal, i.e.,
$q_{t}^{*}=\pi_{t},\ \forall t$ (47)
Case (ii) Otherwise, we have to increase the values of $\{q_{t}\}$ to satisfy the DP constraint.
@@
$\boldsymbol{w}^{(t)*}=\begin{cases}\boldsymbol{w}_{\text{ZF}}^{(t)},&\text{if }\sum_{i=0}^{T-1}\frac{1}{\pi_{i}^{2}}\leq A\\ \frac{q_{t}^{*}}{\pi_{t}}\boldsymbol{w}_{\text{ZF}}^{(t)},&\text{otherwise}\end{cases},\ \ \forall t,$ (51)
@@
Remark IV.1 (Overall approach for solving optimization problem)
The objective of $(\widetilde{\text{P}}\text{0})$, $\sum_{t=0}^{T-1}\|\boldsymbol{w}^{(t)}\|^{2}$, depends solely on the norms of the combining vectors $\{\boldsymbol{w}^{(t)}\}$, not their directions. This enables reformulating $(\widetilde{\text{P}}\text{0})$ as an optimization over the norms $\{q_{t}\}_{t=0}^{T-1}$, where $q_{t}=\|\boldsymbol{w}^{(t)}\|$, subject to feasibility constraints derived from the power constraint (33b) and the DP constraint (33c). For each $t$, the minimum feasible $q_{t}$ is the norm $\pi_{t}\triangleq\|\boldsymbol{w}_{\text{ZF}}^{(t)}\|$ of the ZF solution to $(\text{P1})$ in (36). Thus, any valid $q_{t}$ must satisfy $q_{t}\geq\pi_{t}$. This reduces $(\widetilde{\text{P}}\text{0})$ to $(\widetilde{\text{P}}\text{0}^{\prime\prime})$, which optimizes $q_{t}$} subject to $q_{t}\geq\pi_{t}$ for all $t$ and the DP constraint. Once the optimal norm $q_{t}^{*}$ is found, we construct an optimal solution via scaling, i.e., $\boldsymbol{w}^{(t)*}=(q_{t}^{*}/\pi_{t})\boldsymbol{w}_{\text{ZF}}^{(t)}$, which satisfies the power constraint (33b) due to homogeneity.
@@
  • The ZF solution $\{\boldsymbol{w}^{(t)}_{\text{ZF}}\}_{t=0}^{T-1}$ is optimal if $\sum_{t=0}^{T}1\mathord{\left/\right.\kern-1.2pt}\pi_{t}^{2}\leq A$, i.e., the maximum signal to noise ratio (SNR) satisfies the following inequality
@@
  • Otherwise, the scaled ZF solution $\{(q_{t}^{*}/\pi_{t})\boldsymbol{w}_{\text{ZF}}^{(t)}\}_{t=0}^{T-1}$ is optimal, with $q_{t}^{*}$ obtained by (46).
Remark IV.3 (The property of privacy-for-free is also holds in multi-user SIMO system)
Our results enable a privacy-for-free mechanism, countering the claims in [26] that zero-artificial-noise is not always possible in DP-guaranteed multi-antenna AirFL systems. Specifically, our design eliminates artificial noise in such systems. In low-SNR regimes meeting the condition (52), the DP-aware solution matches ZF, achieving privacy guarantees without performance loss.
Furthermore, it is worth noting that the core findings of our work exhibit a high degree of generality. While related works, such as [26], may adopt different system models—e.g., variations in the specific form of the DP constraint (arising from different privacy quantification methods or threat models) and the power constraint (due to different communication models)—the fundamental procedure for deriving the optimal solution and the structure of the solution itself remain unaffected. Consequently, the key insights of our paper, including the privacy-for-free mechanism and the characterization of the utility-privacy trade-off, are preserved even when adapting our framework to these variations.

### v3

@@ IV-B Problem Formulation
$(\text{P0}):$ $\min_{\{\boldsymbol{w}^{(t)}\}_{t=0}^{T-1},\ \{s_{i}^{(t)}\}_{t=0}^{T-1}}\ \sum_{t=0}^{T-1}\left(\Lambda^{(t)}+\|\boldsymbol{w}^{(t)}\|^{2}d\sigma^{2}\right)$
$\text{ s.t. }\quad c^{2}(s_{i}^{(t)})^{2}\leq dP,\ \forall i\in\mathcal{I}^{(t)},\ \forall t,$ (31a)
$\epsilon\geq\sqrt{\frac{(2c_{\delta}+8)\log(1/\delta)rc^{2}}{\sigma^{2}}\min\left\{\sum_{t=0}^{T-1}\phi_{t},\Phi\right\}}.$ (31b)
Given any $\{\boldsymbol{w}^{(t)}\}_{t=0}^{T-1}$, the optimal $s_{i}^{(t)}$ is given by [40, 38], $s_{i}^{(t)}=1\mathord{\left/\right.\kern-1.2pt}(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}$, which yields $\Lambda^{(t)}=0$, and $\phi_{t}=1/\|\boldsymbol{w}^{(t)}\|^{2}$. By substituting $s_{i}^{(t)}$, $i\in\mathcal{I}^{(t)}$, $\forall t$, we recast problem (P0) into the following equivalent problem.
$(\text{P0}^{\prime}):$ $\min_{\{\boldsymbol{w}^{(t)}\}_{t=0}^{T-1}}\ \sum_{t=0}^{T-1}\|\boldsymbol{w}^{(t)}\|^{2}$
$\text{ s.t. }\quad\left|(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}\right|\geq\frac{c}{dP},\ \forall i\in\mathcal{I}^{(t)},\ \forall t,$ (32a)
$\epsilon\geq\sqrt{\frac{(2c_{\delta}+8)\log(1/\delta)rc^{2}}{\sigma^{2}}\min\left\{\sum_{t=0}^{T-1}\frac{1}{\|\boldsymbol{w}^{(t)}\|^{2}},\Phi\right\}}.$ (32b)
For illustration, we solve a simplified version of $(\text{P0}^{\prime})$ assuming an unbounded parameter domain (i.e., $D\to\infty$), yielding $\min\{\sum_{t}1/\|\boldsymbol{w}^{(t)}\|^{2},\Phi\}=\sum_{t}1/\|\boldsymbol{w}^{(t)}\|^{2}$. Note that the general case with finite $D$ is also tractable and discussed in Appendix D. In this case, $(\text{P0}^{\prime})$ is simplified as:
$(\widetilde{\text{P}}\text{0}):\$$\min_{\{\boldsymbol{w}^{(t)}\}_{t=0}^{T-1}}\quad\sum_{t=0}^{T-1}\|\boldsymbol{w}^{(t)}\|^{2}$
$\text{s.t.}\quad|(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}|\geq c\mathord{\left/\right.\kern-1.2pt}\sqrt{dP},\ \forall i\in\mathcal{I}^{(t)},\ \forall t,$ (33a)
$\quad\ \ \ \ \sum_{t=0}^{T-1}\frac{1}{\|\boldsymbol{w}^{(t)}\|^{2}}\leq A,$ (33b)
@@
Lemma IV.2
Without the DP constraint (33b), $(\widetilde{\text{P}}\text{0})$ can be decoupled into $T$ quadratically constrained quadratic programs (QCQPs), each for one $t$ as follows:
$(\text{P1}):\$$\min_{\boldsymbol{w}^{(t)}}\quad\|\boldsymbol{w}^{(t)}\|^{2}$
$\text{s.t.}\quad|(\boldsymbol{w}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}|\geq c\mathord{\left/\right.\kern-1.2pt}\sqrt{dP},\ \forall i\in\mathcal{I}^{(t)},\ \forall t.$ (35a)
Generally, each sub-problem can be individually addressed by using the semi-definite programming (SDP) methods [20], resulting in the optimal solution to problem $(\text{P1})$, denoted by $\{\boldsymbol{w}_{0}^{(t)}\}$.
Proposition IV.2
@@
$(\widetilde{\text{P}}\text{0}^{\prime}):\$$\min_{\{\boldsymbol{w}^{(t)}\}_{t=0}^{T-1}}\quad\sum_{t=0}^{T-1}\|\boldsymbol{w}^{(t)}\|^{2}$
$\text{ s.t. }\quad\|\boldsymbol{w}^{(t)}\|\geq\|\boldsymbol{w}_{0}^{(t)}\|>0,\ \forall t,$ (36a)
$\quad\ \ \ \ \sum_{t=0}^{T-1}\frac{1}{\|\boldsymbol{w}^{(t)}\|^{2}}\leq A,$ (36b)
where $\boldsymbol{w}^{(t)}_{0}$ denotes the optimal solution to problem $(\text{P1})$.
Proof: To prove this result, we first establish the optimal solution in the absence of the DP constraint (33b), denoted by $\{\boldsymbol{w}_{0}^{(t)}\}$. Denote the optimal solutions to $(\widetilde{\text{P}}\text{0})$ and $(\widetilde{\text{P}}\text{0}^{\prime})$ as $\{\hat{\boldsymbol{w}}^{(t)}\}$ and $\{\widetilde{\boldsymbol{w}}^{(t)}\}$, respectively. We note that the solution $\{\hat{\boldsymbol{w}}^{(t)}\}$ also satisfies the constraint of (36a) and (36b) by Lemma IV.2, making it a feasible solution for $(\widetilde{\text{P}}\text{0}^{\prime})$. We then have
@@
In addition, for the solution $\{\frac{\|\widetilde{\boldsymbol{w}}^{(t)}\|}{\|\boldsymbol{w}^{(t)}_{0}\|}\boldsymbol{w}_{0}^{(t)}\}_{t=0}^{T-1}$, we have
$|\left(\frac{\|\widetilde{\boldsymbol{w}}^{(t)}\|}{\|\boldsymbol{w}^{(t)}_{0}\|}\boldsymbol{w}_{0}^{(t)}\right)^{H}\boldsymbol{h}_{i}^{(t)}|$ $\geq|(\boldsymbol{w}_{0}^{(t)})^{H}\boldsymbol{h}_{i}^{(t)}|$ (38)
@@
$\sum_{t=0}^{T-1}\frac{1}{\|\frac{\|\widetilde{\boldsymbol{w}}^{(t)}\|}{\|\boldsymbol{w}^{(t)}_{0}\|}\boldsymbol{w}_{0}^{(t)}\|^{2}}\leq\sum_{t=0}^{T-1}\frac{1}{\|\widetilde{\boldsymbol{w}}^{(t)}\|^{2}}\leq A,$ (39)
which means that $\bigl\{\frac{\|\widetilde{\boldsymbol{w}}^{(t)}\|}{\|\boldsymbol{w}^{(t)}_{0}\|}\boldsymbol{w}_{0}^{(t)}\bigr\}_{t=0}^{T-1}$ is a feasible solution for $(\widetilde{\text{P}}\text{0})$. Further, we have
$\sum_{t=0}^{T-1}\|\hat{\boldsymbol{w}}^{(t)}\|^{2}\leq\sum_{t=0}^{T-1}\|\frac{\|\widetilde{\boldsymbol{w}}^{(t)}\|}{\|\boldsymbol{w}^{(t)}_{0}\|}\boldsymbol{w}_{0}^{(t)}\|^{2}\leq\sum_{t=0}^{T-1}\|\widetilde{\boldsymbol{w}}^{(t)}\|^{2}.$ (40)
@@
IV-C Optimal Solution to $(\widetilde{\text{P}}\text{0}^{\prime})$
In this subsection, we proceed to derive the optimal solution of problem $(\widetilde{\text{P}}\text{0}^{\prime})$. Let $\pi_{t}=\|\boldsymbol{w}_{0}^{(t)}\|$ and define $q_{t}=\|\boldsymbol{w}^{(t)}\|$. Then, we can rewrite $(\widetilde{\text{P}}\text{0}^{\prime})$ as
$(\widetilde{\text{P}}\text{0}^{\prime\prime}):\$$\min_{\{q_{t}\}_{t=0}^{T-1}}\quad\sum_{t=1}^{T-1}q_{t}^{2}$
$\text{ s.t. }\quad q_{t}\geq\pi_{t}>0,\ \forall t,$ (41a)
$\quad\ \ \ \ \ \ \ \sum_{t=0}^{T-1}\frac{1}{q_{t}^{2}}\leq A,$ (41b)
@@ IV-C Optimal Solution to $(\widetilde{\text{P}}\text{0}^{\prime})$
where $\mu\geq 0$ and $\nu_{t}\geq 0$ are the Lagrangian multipliers associated with constraints (41a) and (41b), respectively. Further, the KKT conditions are given as
@@
where $\mu^{*}$ denotes $\mu$’s optimal value.
@@
Case (i): If $\sum_{t=0}^{T-1}1/\pi_{t}^{2}\leq A$, the solution $\{\pi_{t}\}$ satisfies privacy and is optimal, i.e.,
$q_{t}^{*}=\pi_{t},\ \forall t.$ (45)
Case (ii): Otherwise, we have to increase the values of $\{q_{t}\}$ to satisfy the DP constraint.
@@
$\boldsymbol{w}^{(t)*}=\begin{cases}\boldsymbol{w}_{0}^{(t)},&\text{if }\sum_{i=0}^{T-1}\frac{1}{\pi_{i}^{2}}\leq A\\ \frac{q_{t}^{*}}{\pi_{t}}\boldsymbol{w}_{0}^{(t)},&\text{otherwise}\end{cases},\ \ \forall t,$ (49)
@@
Remark IV.1
The objective of $(\widetilde{\text{P}}\text{0})$, $\sum_{t=0}^{T-1}\|\boldsymbol{w}^{(t)}\|^{2}$, depends solely on the norms of the combining vectors $\{\boldsymbol{w}^{(t)}\}$, not their directions. This enables reformulating $(\widetilde{\text{P}}\text{0})$ as an optimization over the norms $\{q_{t}\}_{t=0}^{T-1}$, where $q_{t}=\|\boldsymbol{w}^{(t)}\|$, subject to feasibility constraints derived from the power constraint (33a) and the DP constraint (33b). For each $t$, the minimum feasible $q_{t}$ is the norm $\pi_{t}\triangleq\|\boldsymbol{w}_{0}^{(t)}\|$ of the optimal solution to $(\text{P1})$. Thus, any valid $q_{t}$ must satisfy $q_{t}\geq\pi_{t}$. This reduces $(\widetilde{\text{P}}\text{0})$ to $(\widetilde{\text{P}}\text{0}^{\prime\prime})$, which optimizes $q_{t}$} subject to $q_{t}\geq\pi_{t}$ for all $t$ and the DP constraint. Once the optimal norm $q_{t}^{*}$ is found, we construct an optimal solution via scaling, i.e., $\boldsymbol{w}^{(t)*}=(q_{t}^{*}/\pi_{t})\boldsymbol{w}_{0}^{(t)}$, which satisfies the power constraint (33a) due to homogeneity.
@@
  • The solution $\{\boldsymbol{w}^{(t)}_{0}\}_{t=0}^{T-1}$ is optimal if $\sum_{t=0}^{T}1\mathord{\left/\right.\kern-1.2pt}\pi_{t}^{2}\leq A$. In particular, under the special case where $m\geq rn$ and the channel coefficients are i.i.d., $\{\pi_{t}\}_{t=0}^{T-1}$ can be expressed as [7]:
    $\pi_{t}=\|\frac{c}{\sqrt{dP}}\boldsymbol{H}^{(t)}\left((\boldsymbol{H}^{(t)})^{H}\boldsymbol{H}^{(t)}\right)^{-1}\boldsymbol{u}\|,$ (50)
    where $\boldsymbol{H}^{(t)}=[\boldsymbol{h}_{1}^{(t)},\cdots,\boldsymbol{h}_{rn}^{(t)}]\in\mathbb{C}^{m\times rn}$, and $\boldsymbol{u}\in\mathbb{C}^{rn}$ denotes the vector whose components all have unit modulus. Then, the maximum signal to noise ratio (SNR) satisfies the following inequality
@@
  • Otherwise, the scaled solution $\{(q_{t}^{*}/\pi_{t})\boldsymbol{w}_{0}^{(t)}\}_{t=0}^{T-1}$ is optimal, with $q_{t}^{*}$ obtained by (44).
Remark IV.3
Our results enable a privacy-for-free mechanism, countering the claims in [18] that zero-artificial-noise is not always possible in DP-guaranteed multi-antenna AirFL systems. Specifically, our design eliminates artificial noise in such systems. In low-SNR regimes meeting the condition (51), the DP-aware solution matches the optimal solution to $(\text{P1})$, achieving privacy guarantees without performance loss.
Furthermore, it is worth noting that the core findings of our work exhibit a high degree of generality. While related works, such as [18], may adopt different system models—e.g., variations in the specific form of the DP constraint (arising from different privacy quantification methods or threat models) and the power constraint (due to different communication models)—the fundamental procedure for deriving the optimal solution and the structure of the solution itself remain unaffected. Consequently, the key insights of our paper, including the privacy-for-free mechanism and the characterization of the convergence-privacy trade-off, are preserved even when adapting our framework to these variations.
Figure 3: The evolution of the (a) testing accuracy and (b) training loss versus the global communication rounds $T$ with different DP budgets $\tilde{\epsilon}$ under the i.i.d. setting.
Figure 4: The evolution of the (a) testing accuracy and (b) training loss versus the global communication rounds $T$ with different DP budgets $\tilde{\epsilon}$ under the non-i.i.d. setting.